# OpenOrbitalOptimizer—a reusable open source library for self-consistent field calculations

### v1

@@ 2 Theory
where the Fock and overlap matrices are complex for $\boldsymbol{k}\neq\boldsymbol{0}$; note here that $\boldsymbol{k}$ and $\sigma$ label symmetry blocks. Since SCF acceleration in periodic calculations follows closely the molecular approach,69 and eq. 9 is still of the same form as eq. 2, the library should also support periodic calculations without changes.
@@ 2.1.1 Pulay DIIS
in order to speed up the slow convergence of the iterative Roothaan procedure. The parameter $m$ in eq. 15 is the length of the history. The length of the history is limited by default to $m=10$ in OpenOrbitalOptimizer, following the default in Psi4.54 Note that there is a wide variation for the default value in various programs, e.g. ORCA70 employs $m=5$, Q-Chem71 $m=15$, and Gaussian72 employs $m=20$ by default.
@@
For large history sizes, the error vectors can become linearly dependent. Hamilton and Pulay 73 suggested multiplying the diagonal elements of eq. 20 by a factor $1+d$ where $d=0.02$ is a damping factor.
As the last row and column of eq. 20 is always of the order of unity, the numerical conditioning of the linear group of equations can be improved by scaling the errors given by eq. 19 such that the most recent diagonal element becomes unity.74 In our implementation, the matrix is normalized with respect to the smallest diagonal element $B_{ii}$, instead; this should have no effect on the numerical behavior.
In addition, there is literature suggesting that limiting the matrices included in DIIS extrapolation may help in some cases. We use the simple criterion suggested by Chupin et al. 75 for restarting DIIS: only those vectors $i$ of the current history are included in the extrapolation that satisfy
@@ 2.1.2 EDIIS and ADIIS
DIIS can be viewed as a quasi-Newton method,76, 75 and it generally leads to fast convergence when the orbitals are close to convergence, i.e., when the orbital gradient is small. However, this also means that DIIS can only be trusted to work when one is close to a minimum; otherwise, the method can also easily converge onto a saddle point. If the initial guess is poor, other methods must therefore be used instead.
Energy DIIS22 (EDIIS) and augmented DIIS23 (ADIIS) are based on proxy energy functionals, which allow the methods to work far from the optimum. The two methods are equivalent in the case of Hartree–Fock theory,77 but differ in the case of density-functional theory1, 2 (DFT) for which the energy is not a quadratic function of the density matrix.
@@
and such implementations appear to be widely used, e.g. in ERKALE,50 Psi4,54 PySCF,55 and Q-Chem.71 However, we have found that this technique can sometimes fail in the line search. This issue is easily understood in the case where the minimum features $c_{i}=0$ for some $c_{i}$: while the line searches will try to make the ratios of the components of $\boldsymbol{x}$ larger, the resulting $\boldsymbol{c}$ parameters are invariant to the scaling $\boldsymbol{x}\to\lambda\boldsymbol{x}$, which means that the optimization becomes slowly convergent.
Since the problem is usually rather low-dimensional, being fully determined by the length of the history, we will use a simple method inspired by Powell’s method,78 instead, which also appears to have been used as a fallback by Kudin et al. 22.
@@ 2.1.3 Hybrid scheme of EDIIS/ADIIS followed by DIIS
Because EDIIS and ADIIS work well far from the optimum, while DIIS works well close to it, we follow Garza and Scuseria 77 and interpolate between the two sets of weights
@@
[…]
@@
Again following Garza and Scuseria 77, pure EDIIS/ADIIS is used if the error of the last cycle is 10% greater than the minimum error.
To choose between EDIIS and ADIIS in a given iteration of eq. 31, we employ the minimal error sampling algorithm (MESA) of Garcia Chavez 80: we compute both the EDIIS and ADIIS extrapolations, and pick the one that yields the larger projection onto the current density matrix.
@@ 2.1.4 History cleanup
Even though ADIIS and EDIIS are interpolation methods that are based on solid mathematical models, using faraway points can still lead to degraded performance in DFT calculations: our exploratory calculations showed that when the orbitals undergo large changes in occupations, it is beneficial to clear the history. In our implementation, following the idea of Chupin et al. 75 discussed above in eq. 21, we keep only those density matrices in the history that satisfy
@@ 3 Design Choices
Arguably the most important design choice in any project is the license. In order to promote the reusability of the library and avoid the need for duplicate solutions,36 we chose to license OpenOrbitalOptimizer under the Mozilla Public License that is also used in Libxc,82 which is nowadays used by over 40 programs including also several commercial programs.
@@
Moreover, as the main tasks of the library have to do with linear algebra, the library is developed on top of the Armadillo library,83 which offers a convenient syntax for matrix and vector operations, as well as the required support for templated datatypes that are used throughout OpenOrbitalOptimizer. The use of Armadillo within the library does not restrict the use of the library in other codes, since the matrix data is easy to convert to other formats in the library interface.
@@ 4 Computational Details
Although OpenOrbitalOptimizer has already been interfaced with several programs, such as HelFEM51, 52, 84 and Psi4,54 we exemplify the library with nuclear-electronic orbital62, 63, 64 (NEO) Hartree–Fock (HF) calculations that we implemented in a plugin to the ERKALE program.50, 85
@@
As electronic basis sets, we consider the polarization consistent pc-n and aug-pc-n basis sets of Jensen.86, 87 Following our recent work in ref. 88, we uncontract the electronic basis set on the quantum protons, as this leads to significant improvements in the precision of the calculation; this procedure is denoted as uncHq-. The protonic basis set is PB4-F189 in all calculations. All basis sets are taken from the Basis Set Exchange.90
We use density fitting in all calculations with autogenerated auxiliary basis sets. As recently reviewed by Pedersen et al. 91, there is a close connection between density fitting and Cholesky decompositions. Liu et al. 92 showed that the Cholesky decomposition yields good results in NEO calculations when the decomposition is carried out for the union of the protonic and electronic basis functions; we thus generated the auxiliary basis set from pivoted Cholesky decompositions for the union of the protonic and electronic basis functions following our recent work.93, 94 We then used the same auxiliary basis set for both the electrons and the protons.
We assess the accuracy of our density-fitted NEO-HF implementation against calculations performed with exact integrals in Q-Chem.71 The two-electron integral threshold was tightened to $\epsilon=10^{-12}$ in Q-Chem. The basis set linear dependence thresholds were set to $10^{-5}$ in both codes. The Coulomb overlap matrix in the density fitting approach was inverted using the pivoted Cholesky technique of ref. 14, using a $10^{-8}$ threshold to choose the pivot functions, and a $10^{-7}$ linear dependence threshold for the canonical orthonormalization step.
@@
Instead, we found that a minimal 1s protonic basis guess with exponent $16\sqrt{2}\approx 22.627$ adopted from the 8s8p8d even-tempered basis of Yang et al. 95 leads to qualitatively correct electronic and protonic orbitals, as calculations in the larger basis sets with this read-in guess converge to the DIIS region in just a few iterations. Performing this initial calculation has practically no overhead, since there are no protonic degrees of freedom to optimize in the minimal basis.
We note that Liu et al. 65 described using a guess where the proton occupies the tightest protonic basis function for each quantum nucleus—which we assume to refer to an s function. Our guess is thus somewhat similar to that of Liu et al., but our guess is independent of the protonic basis used in the actual calculation. Note that the tightest basis function in the protonic 8s8p8d basis of Yang et al. 95 has an exponent of 32.0, while the tightest s function in PB4-F189 has an exponent 28.950.
@@
Due to the differences in Fock matrix constructions, the simultaneous and stepwise approaches can have very different costs per iteration, as illustrated by table 1. Even though the simultaneous optimization has been found to converge faster,65 the electronic basis set is much larger than the protonic one in typical calculations; this is especially the case when one is interested in going towards the complete basis set limit, see Khan and Tonner-Zech 96 and our recent work in ref. 88. Even though the protonic problem tends to be more difficult to converge, requiring many more iterations than the electronic one, the protonic iterations are usually extremely cheap in comparison to the electronic ones.
@@ 5 Results
Calculations were furthermore performed with auxiliary basis sets of three sizes: the “small”, “large”, and “verylarge” auxiliary basis sets generated with the procedure of ref. 94. As previous experience shows that at least a triple-$\zeta$ orbital basis is required for generating reliable auxiliary basis sets,93 data presented for the double-$\zeta$ pc-1 and aug-pc-1 orbital basis sets employ the auxiliary basis set of the corresponding triple-$\zeta$ orbital basis set.
@@
Table 2: Number of Fock matrix builds needed to converge the NEO-HF calculations with the “large” auxiliary basis: $N_{\textrm{sim}}$ is the number of Fock builds in the simultaneous electronic-protonic SCF, while $N_{\text{elec}}$ and $N_{\text{prot}}$ are the total numbers of electronic and protonic Fock builds in the stepwise SCF procedure (c.f. table 1). $N_{\textrm{ep/pe}}$ shows the number of times the electron-proton and proton-electron coupling terms were calculated, while $\Delta E=E(\text{simultaneous})-E(\text{separate})$ shows the difference in the final total energies.
@@ 6 Summary and Conclusions
We have presented OpenOrbitalOptimizer,45 a reusable open-source header-only templated C++ library for orbital optimization and self-consistent field (SCF) calculations. OpenOrbitalOptimizer aims to either fully eliminate existing disparate SCF implementations in various programs, or alternatively to supplement their functionalities by offering access to consistent implementations of various SCF methodologies.
OpenOrbitalOptimizer offers a unified interface for solving coupled SCF equations of an arbitrary number of types of particles, with an arbitrary number of symmetries. OpenOrbitalOptimizer has already been interfaced with HelFEM,84, 51, 52, 60, 53, 57 ERKALE,50, 85 and Psi4,54 and we hope to see many more interfaces with other programs.
@@
In future work, we aim to extend the functionality of OpenOrbitalOptimizer for calculations on systems with degeneracies at the Fermi level, in which case the Aufbau theorem does not hold and the occupations of the degenerate orbitals must instead be explicitly optimized. Although our main interest in this problem arises from atomic calculations,98, 10, 53, 57, 59 degenerate occupations can sometimes be also found in molecules,99, 100, 101 and are likely the main challenge in large-scale black-box applications of DFT. A reusable implementation of such an approach in OpenOrbitalOptimizer would then allow exploring fractional-occupation solutions in any context, given the flexibility of the present approach.
OpenOrbitalOptimizer is openly available free of charge under the Mozilla Public License, with the main project hosted at https://github.com/susilehtola/OpenOrbitalOptimizer. It runs on Linux, MacOS, and Windows, and will be available through Fedora and conda-forge package managers.
CRediT author statement
Acknowledgments
We dedicate this work to the memory of John F Stanton, who was a positive influence on the community. SL thanks Eric Cancès for many discussions on orbital optimization and the optimal damping algorithm, Mi-Song Dupuy for discussions on the DIIS algorithm, Antoine Levitt for discussions on finite precision issues, his brother Ville Lehtola for discussions on reproducible science, as well as Alejandro Garza and Gustavo Scuseria for confirming the form of the EDIIS+DIIS interpolation function. SL thanks the Academy of Finland for financial support under project numbers 350282 and 353749. SL also acknowledges the critical support of a French–Finnish Maupertuis researcher short mobility grant funded by the French Institute in Finland, the French Ministry of Higher Education and Research, and the Finnish Society of Sciences and Letters at an early stage of this project.
@@ Acknowledgments
As noted by Kudin et al. 22, the manipulation is straightforward; it suffices to note that one first needs to insert $1=\sum_{j}c_{j}$ to arrive at an expression which looks like $\sum_{ij}A_{ij}P_{i}$, where $\boldsymbol{A}$ is an antisymmetric quantity, $A_{ij}=-A_{ji}$. In the second step, splitting the sum in two copies and interchanging the indices $i\leftrightarrow j$ in the other term one, gets $\sum_{ij}A_{ij}P_{i}=\frac{1}{2}\sum_{ij}A_{ij}P_{i}+\frac{1}{2}\sum_{ji}A_{ji}P_{j}=\frac{1}{2}\sum_{ij}A_{ij}(P_{i}-P_{j})$. These two steps are sufficient to manipulate eq. 46 into the form
@@
$-\frac{1}{4}\sum_{ij}c_{i}c_{j}\sum_{\sigma}\text{Tr }\left[\boldsymbol{F}_{\sigma}\left(\boldsymbol{P}_{\sigma}^{(j)}\right)-\boldsymbol{F}_{\sigma}\left(\boldsymbol{P}_{\sigma}^{(i)}\right)\right]\times$

### v2

@@ 2 Methods
where the Fock and overlap matrices are in general complex for $\boldsymbol{k}\neq\boldsymbol{0}$; note here that $\boldsymbol{k}$ and $\sigma$ label symmetry blocks. Since SCF acceleration in periodic calculations follows closely the molecular approach,69 and eq. 9 is still of the same form as eq. 2, the library should also support periodic calculations without changes. If necessary, the code is straightforward to amend to allow k-point weighting in the convergence acceleration.70, 69
Fully relativistic four-component calculations offer another example of the need of complex-valued SCF, relativistic effects being famously important for the chemistry of heavy elements.71, 72 The main complication in four-component calculations is the Dirac sea of negative-energy solutions, which correspond to positronic solutions.73 One should only populate the positive-energy orbitals that correspond to electrons, and this is achieved by skipping over the negative-energy solutions when determining the occupations of the four-component spinors.73 The number of negative-energy solutions is given by the number of (linearly independent) small-component basis functions in the calculation, which is thus known beforehand. Although this feature does not exist at present in OpenOrbitalOptimizer, the required change is thus trivial.
Note that relativistic effects are often described within approaches that do not feature the Dirac sea, the negative energy states being projected out by design. Popular approaches in present-day quantum chemistry include the use of effective core potentials,74 and the use of the exact two-component (X2C) method.75, 76, 77, 78 In contrast to the four-component calculations discussed in the previous paragraph, these methods do not cause any extra complications, as the SCF is carried out for these methods exactly in the same way as in the non-relativistic case.
@@ 2.1.1 Pulay DIIS
in order to speed up the slow convergence of the iterative Roothaan procedure. The parameter $m$ in eq. 15 is the length of the history. The length of the history is limited by default to $m=10$ in OpenOrbitalOptimizer, following the default in Psi4.54 Note that there is a wide variation for the default value in various programs, e.g. ORCA79 employs $m=5$, Q-Chem80 $m=15$, and Gaussian81 employs $m=20$ by default.
@@
For large history sizes, the error vectors can become linearly dependent. Hamilton and Pulay 82 suggested multiplying the diagonal elements of eq. 20 by a factor $1+d$ where $d=0.02$ is a damping factor.
As the last row and column of eq. 20 is always of the order of unity, the numerical conditioning of the linear group of equations can be improved by scaling the errors given by eq. 19 such that the most recent diagonal element becomes unity.83 In our implementation, the matrix is normalized with respect to the smallest diagonal element $B_{ii}$, instead; this should have no effect on the numerical behavior.
In addition, there is literature suggesting that limiting the matrices included in DIIS extrapolation may help in some cases. We use the simple criterion suggested by Chupin et al. 84 for restarting DIIS: only those vectors $i$ of the current history are included in the extrapolation that satisfy
@@ 2.1.2 EDIIS and ADIIS
DIIS can be viewed as a quasi-Newton method,85, 84 and it generally leads to fast convergence when the orbitals are close to convergence, i.e., when the orbital gradient is small. However, this also means that DIIS can only be trusted to work when one is close to a minimum; otherwise, the method can also easily converge onto a saddle point. If the initial guess is poor, other methods must therefore be used instead.
Energy DIIS22 (EDIIS) and augmented DIIS23 (ADIIS) are based on proxy energy functionals, which allow the methods to work far from the optimum. The two methods are equivalent in the case of Hartree–Fock theory,86 but differ in the case of density-functional theory1, 2 (DFT) for which the energy is not a quadratic function of the density matrix.
@@
and such implementations appear to be widely used, e.g. in ERKALE,50 Psi4,54 PySCF,55 and Q-Chem.80 However, this is a bad idea since the minimization problem is no longer convex in $\boldsymbol{x}$. For example, we have found that this technique sometimes fails in the line search. This issue is easily understood in the case where the minimum features $c_{i}=0$ for some $c_{i}$: while the line searches will try to make the ratios of the components of $\boldsymbol{x}$ larger, the resulting $\boldsymbol{c}$ parameters are invariant to the scaling $\boldsymbol{x}\to\lambda\boldsymbol{x}$, which means that the optimization becomes slowly convergent.
Since the problem is usually rather low-dimensional, being fully determined by the length of the history, we will use a simple method inspired by Powell’s method,87 instead, which also appears to have been used as a fallback by Kudin et al. 22.
@@ 2.1.3 Hybrid scheme of EDIIS/ADIIS followed by DIIS
Because EDIIS and ADIIS work well far from the optimum, while DIIS works well close to it, we follow Garza and Scuseria 86 and interpolate between the two sets of weights
@@
Although the form eq. 32 is not obvious in the equation at the bottom of page 3 of ref. 86, it appears that Garza and Scuseria 86 indeed used
@@
We note that Garza and Scuseria 86 suggested using pure EDIIS/ADIIS if the error of the last cycle is 10% greater than the minimum error. We note that the error can also increase even when the energy is lowered, and that we found this criterion to sometimes lead to convergence to local minima. For this reason, we decided not to implement this feature in OpenOrbitalOptimizer.
To choose between EDIIS and ADIIS in a given iteration of eq. 31, we employ the minimal error sampling algorithm (MESA) of Garcia Chavez 88: we compute both the EDIIS and ADIIS extrapolations, and pick the one that results in the least change to the density matrix as measured by the Frobenius norm $\left\|\boldsymbol{P}_{\text{current}}-\boldsymbol{P}_{\text{new}}\right\|_{F}$.
@@ 2.1.4 History cleanup
Even though ADIIS and EDIIS are interpolation methods that are based on solid mathematical models, using faraway points can still lead to degraded performance in DFT calculations: our exploratory calculations showed that when the orbitals undergo large changes in occupations, it is beneficial to clear the history. In our implementation, following the idea of Chupin et al. 84 discussed above in eq. 21, we keep only those density matrices in the history that satisfy
@@ 3 Design Choices
Arguably the most important design choice in any project is the license. In order to promote the reusability of the library and avoid the need for duplicate solutions,36 we chose to license OpenOrbitalOptimizer under the Mozilla Public License that is also used in Libxc,90 which is nowadays used by over 40 programs including also several commercial programs.
@@
Moreover, as the main tasks of the library have to do with linear algebra, the library is developed on top of the Armadillo library,91 which offers a convenient syntax for matrix and vector operations, as well as the required support for templated datatypes that are used throughout OpenOrbitalOptimizer. The use of Armadillo within the library does not restrict the use of the library in other codes, since the matrix data is easy to convert to other formats in the library interface.
@@
This procedure has the potential pitfall in that it can get stuck. We decided to resolve this issue by triggering an alternative strategy in the case $N/2$ consecutive (A/E)DIIS iterations have failed to result in a decrease of the total energy, $N$ being the maximum allowable history length: our tit-for-tat strategy is to carry out the next $N/2$ iterations with ODA in the hopes to make sure that the solution gets sufficiently far away from the stagnant region for (A/E)DIIS to recover its usefulness.
@@ 4 Computational Details
Although OpenOrbitalOptimizer has already been interfaced with several programs, such as HelFEM51, 52, 92 and Psi4,54 we exemplify the library with nuclear-electronic orbital62, 63, 64 (NEO) Hartree–Fock (HF) calculations that we implemented in a plugin to the ERKALE program.50, 93
@@
As electronic basis sets, we consider the polarization consistent pc-n and aug-pc-n basis sets of Jensen.94, 95 Following our recent work in ref. 96, we uncontract the electronic basis set on the quantum protons, as this leads to significant improvements in the precision of the calculation; this procedure is denoted as uncHq-. The protonic basis set is PB4-F197 in all calculations. All basis sets are taken from the Basis Set Exchange.98
We use density fitting in all calculations with autogenerated auxiliary basis sets. As recently reviewed by Pedersen et al. 99, there is a close connection between density fitting and Cholesky decompositions. Liu et al. 100 showed that the Cholesky decomposition yields good results in NEO calculations when the decomposition is carried out for the union of the protonic and electronic basis functions; we thus generated the auxiliary basis set from pivoted Cholesky decompositions for the union of the protonic and electronic basis functions following our recent work.101, 102 We then used the same auxiliary basis set for both the electrons and the protons.
We assess the accuracy of our density-fitted NEO-HF implementation against calculations performed with exact integrals in Q-Chem.80 The two-electron integral threshold was tightened to $\epsilon=10^{-12}$ in Q-Chem. The basis set linear dependence thresholds were set to $10^{-5}$ in both codes. The Coulomb overlap matrix in the density fitting approach was inverted using the pivoted Cholesky technique of ref. 14, using a $10^{-8}$ threshold to choose the pivot functions, and a $10^{-7}$ linear dependence threshold for the canonical orthonormalization step.
@@
Instead, we found that a minimal 1s protonic basis guess with exponent $16\sqrt{2}\approx 22.627$ adopted from the 8s8p8d even-tempered basis of Yang et al. 103 leads to qualitatively correct electronic and protonic orbitals, as calculations in the larger basis sets with this read-in guess converge to the DIIS region in just a few iterations. Performing this initial calculation has practically no overhead, since there are no protonic degrees of freedom to optimize in the minimal basis.
We note that Liu et al. 65 described using a guess where the proton occupies the tightest protonic basis function for each quantum nucleus—which we assume to refer to an s function. Our guess is thus somewhat similar to that of Liu et al., but our guess is independent of the protonic basis used in the actual calculation. Note that the tightest basis function in the protonic 8s8p8d basis of Yang et al. 103 has an exponent of 32.0, while the tightest s function in PB4-F197 has an exponent 28.950.
@@
Due to the differences in Fock matrix constructions, the simultaneous and stepwise approaches can have very different costs per iteration, as illustrated by table 1. Even though the simultaneous optimization has been found to converge faster,65 the electronic basis set is much larger than the protonic one in typical calculations; this is especially the case when one is interested in going towards the complete basis set limit, see Khan and Tonner-Zech 104 and our recent work in ref. 96. Even though the protonic problem tends to be more difficult to converge, requiring many more iterations than the electronic one, the protonic iterations are usually extremely cheap in comparison to the electronic ones.
@@ 5 Results
Calculations were furthermore performed with auxiliary basis sets of three sizes: the “small”, “large”, and “verylarge” auxiliary basis sets generated with the procedure of ref. 102. As previous experience shows that at least a triple-$\zeta$ orbital basis is required for generating reliable auxiliary basis sets,101 data presented for the double-$\zeta$ pc-1 and aug-pc-1 orbital basis sets employ the auxiliary basis set of the corresponding triple-$\zeta$ orbital basis set.
@@
Table 2: Number of Fock matrix builds needed to converge the NEO-HF calculations with the “large” auxiliary basis for various geometries of the protonated water tetramer: $N_{\textrm{sim}}$ is the number of Fock builds in the simultaneous electronic-protonic SCF, while $N_{\text{elec}}$ and $N_{\text{prot}}$ are the total numbers of electronic and protonic Fock builds in the stepwise SCF procedure (c.f. table 1). $N_{\textrm{ep/pe}}$ shows the number of times the electron-proton and proton-electron coupling terms were calculated, while $\Delta E=E(\text{simultaneous})-E(\text{separate})$ shows the difference in the final total energies.
@@ 6 Summary and Conclusions
We have presented OpenOrbitalOptimizer,45 a reusable open-source header-only templated C++ library for orbital optimization and self-consistent field (SCF) calculations. OpenOrbitalOptimizer aims to either fully eliminate existing disparate SCF implementations in various programs, or alternatively to supplement their functionalities by offering access to consistent implementations of various SCF methodologies. Attesting to the power of modern programming paradigms, including extensive comments, the present implementation is only around 2400 lines of C++ code.
OpenOrbitalOptimizer offers a unified interface for solving coupled SCF equations of an arbitrary number of types of particles, with an arbitrary number of symmetries. OpenOrbitalOptimizer has already been interfaced with HelFEM,92, 51, 52, 60, 53, 57 ERKALE,50, 93 and Psi4,54 and we hope to see many more interfaces with other programs. While none of the currently interfaced codes perform fully relativistic or solid-state calculations, the necessary extensions discussed in the Introduction should be straightforward to make in OpenOrbitalOptimizer.
OpenOrbitalOptimizer is openly available free of charge under the Mozilla Public License, with the main project hosted at https://github.com/susilehtola/OpenOrbitalOptimizer. It runs on Linux, MacOS, and Windows.
@@
In future work, we aim to extend the functionality of OpenOrbitalOptimizer for calculations on systems with degeneracies at the Fermi level, in which case the Aufbau theorem does not hold and the occupations of the degenerate orbitals must instead be explicitly optimized. Although our main interest in this problem arises from atomic calculations,106, 10, 53, 57, 59 degenerate occupations can sometimes be also found in molecules,107, 108, 109 and are likely the main challenge in large-scale black-box applications of DFT. A reusable implementation of such an approach in OpenOrbitalOptimizer would then allow exploring fractional-occupation solutions in any context, given the flexibility of the present approach.
One other potential avenue for future development are interfaces to other languages, such as Python and Fortran, which is easiest to accomplish via a C language interface.36 We hope to be able to accommodate a C language interface in the future. While the C++ library is header-only, the C interface will result in a compiled binary, which may complicate software distribution.
@@
As noted by Kudin et al. 22, the manipulation is straightforward; it suffices to note that one first needs to insert $1=\sum_{j}c_{j}$ to arrive at an expression which looks like $\sum_{ij}A_{ij}P_{i}$, where $\boldsymbol{A}$ is an antisymmetric quantity, $A_{ij}=-A_{ji}$. In the second step, splitting the sum in two copies and interchanging the indices $i\leftrightarrow j$ in the other term, one gets $\sum_{ij}A_{ij}P_{i}=\frac{1}{2}\sum_{ij}A_{ij}P_{i}+\frac{1}{2}\sum_{ji}A_{ji}P_{j}=\frac{1}{2}\sum_{ij}A_{ij}(P_{i}-P_{j})$. These two steps are sufficient to manipulate eq. 46 into the form
@@
$-\frac{1}{4}\sum_{ij}c_{i}c_{j}\sum_{\sigma}\text{Tr }$
$\left[\boldsymbol{F}_{\sigma}\left(\boldsymbol{P}_{\sigma}^{(j)}\right)-\boldsymbol{F}_{\sigma}\left(\boldsymbol{P}_{\sigma}^{(i)}\right)\right]\times$
@@
CRediT author statement
Acknowledgments
We dedicate this work to the memory of John F Stanton, who was a positive influence on the community and a supporter of connecting software. SL thanks Eric Cancès for many discussions on orbital optimization and the optimal damping algorithm, Mi-Song Dupuy for discussions on the DIIS algorithm, Antoine Levitt for discussions on finite precision issues, his brother Ville Lehtola for discussions on reproducible science, as well as Alejandro Garza and Gustavo Scuseria for confirming the form of the EDIIS+DIIS interpolation function. SL thanks the Academy of Finland for financial support under project numbers 350282 and 353749. SL also acknowledges the critical support of a French–Finnish Maupertuis researcher short mobility grant funded by the French Institute in Finland, the French Ministry of Higher Education and Research, and the Finnish Society of Sciences and Letters at an early stage of this project.
{tocentry}[Uncaptioned image]